# Accurate and Diverse LLM Mathematical Reasoning via Automated PRM-Guided GFlowNets

### v2

@@ 3.1 Automated Data Generation via MCTS
17:     if $correct\_count=0$ then
18:        Add $s_{\text{candidate}}$ to the dataset with value 00
@@
We applied this MCTS data generation process (Algorithm 1) using the Qwen2.5-Math model to an initial set of 70,000 mathematical problems sourced from the OpenMathInstruct2 dataset [37].
3.2 Dataset Augmentation via Rollout Reuse and Similarity Grouping
However, directly adding all rollout steps without further processing can introduce inconsistencies. As illustrated in Table 1, semantically similar steps can receive different $MC(s)$ values if evaluated in isolation with a limited number of rollouts ($k=1$, which would be the case if we directly evaluated each step in a rollout independently). This is because evaluating with $k=1$ can lead to noisy and unreliable step value estimations.
Table 1: Example of Similar Steps with Different Values
 Steps value
Max attended 40 college courses in 2 years 0
Max attended 40 college courses 1
  • If all steps within a group originated from rollouts that led to correct final answers, all steps in the group are assigned a value of 1.
  • If all steps within a group originated from rollouts that led to incorrect final answers, all steps in the group are assigned a value of 0.
  • If there is a mixture of correct and incorrect rollouts associated with the steps in a group, all steps in the group are assigned the $MC(s)$ value that was originally estimated for the MCTS step from which these rollouts were generated. In this case, we use the original $MC(s)$ as a reasonable approximation of the value for all similar steps in the group, assuming that steps similar to a step with a known $MC(s)$ are likely to have a similar probability of leading to a correct solution.
3.3 PRM Training
We fine-tuned Qwen2.5-Math [30] as our PRM. We train the PRM to predict the probability of a step leading to a correct solution using Binary Cross-Entropy Loss (BCELoss):
$L=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\right],$
where $y_{i}$ is the true label for the $i$-th step (i.e., the $MC(s)$ value), and $\hat{y}_{i}$ is the predicted probability for the $i$-th step, given by:
@@ 3.3 PRM Training
3.4 PRM Validation
We validate our PRM through three complementary analyses. First, we assess its ability to detect step-level errors by corrupting correct solution steps in systematic ways (e.g., changing numbers, or removing key reasoning components). The PRM consistently assigns lower scores to corrupted steps, demonstrating its sensitivity to reasoning errors.
Second, we evaluate its capability to support diverse solution paths by comparing scores assigned to different valid approaches for the same problem. The PRM assigns comparable scores (within 0.1) to different valid approaches, indicating its ability to recognize multiple correct reasoning paths. Detailed validation results and example cases are provided in subsections 3.4.1 and 3.4.2.
Thirdly, we implemented a guided search strategy to enhance the mathematical reasoning capabilities of LLMs using our trained PRM, similar to the approach in Snell et al. [35]. At each step of generation, the LLM proposes $k$ candidate steps (generated with a temperature of 0.8 for diversity), and the PRM selects the step with the highest score. This step is then appended to the prompt, and the process is repeated until a complete solution is generated.
Figure 2: Accuracy of Falcon models on a subset of MATH Hard using PRM-guided search with varying numbers of proposed steps ($k$). Dashed horizontal lines indicate the baseline accuracy of unguided Falcon3-3B, 7B, and 10B models using prompt-based decoding. Curves represent the accuracy of Falcon3-1B, 3B, and 7B using PRM-guided search with different $k$ values. This demonstrates the effectiveness of PRM-guided search in improving LLM accuracy on challenging mathematical reasoning tasks, especially for smaller models.
Figure 2 demonstrates the effectiveness of this PRM-guided search on a subset of the MATH Hard dataset. The figure compares the accuracy of different Falcon models (1B, 3B, and 7B) using PRM-guided search with varying numbers of proposed steps ($k$) against the baseline accuracy of larger, unguided models using prompt-based decoding.
These results highlight the effectiveness of our PRM in guiding the step-by-step reasoning of LLMs. By selecting the most promising steps based on the PRM’s evaluation, the guided search strategy significantly enhances the accuracy of smaller models, even allowing them to surpass the performance of larger, unguided models on challenging mathematical reasoning tasks. Similar use of PRMs can be found in Wang et al. [39, 40] or other “helper” models such as Preference Process Model in Guan et al. [10]
3.4.1 Error Detection Validation
We systematically corrupted correct solution steps in three ways: Number manipulation, Logic corruption, and Step reordering. The PRM showed consistent score reductions.
3.4.2 Solution Diversity Validation
We evaluated PRM scores on MATH problems with multiple solutions. The PRM assigned comparable scores to different valid reasoning steps, indicating support for multimodal reasoning as can be seen in Figure3.
4 Step-Level GFlowNet Fine-tuning for Diverse Solutions
4.1 Framework definition
Building on Hu et al. [13], we adapt GFlowNets to operate at the reasoning step level rather than the token level. In our framework, a state $s$ represents a partial solution (the question and generated reasoning steps), and an action corresponds to appending an extra step to $s$ to obtain another partial solution $s^{\prime}$. This step-level granularity allows more semantically meaningful control over solution generation.
The forward policy $\pi(s^{\prime}|s)$, parameterized by an LLM, defines the probability of generating step $s^{\prime}$ given state $s$. Following Bengio et al. [2], we incorporate a sink state $s_{f}$ to handle variable-length solutions, with $\pi(s_{f}|s)$ representing the termination probability. The reward $R(s_{n})$ is only defined for states that can transition to $s_{f}$ (terminating states) and is computed as:
$R(s_{n})=\prod_{i=1}^{n}U(s_{i}|s_{i-1}).$ (6)
During generation, we use temperature sampling ($T=0.6$) to promote diversity in the proposed steps. This is particularly important as GFlowNets are designed to sample from the reward-proportional distribution rather than simply maximizing rewards.
4.2 Training Objective and Implementation
We use the modified SubTB loss [21] of Hu et al. [13], adapted for our step-level setting:
Our training procedure maintains a replay buffer $\mathcal{B}$ of size 1000 storing complete trajectories. The replay buffer prioritizes trajectories with higher rewards to accelerate learning. For each question, we:
  1. 1.
    Generate $k$ candidate reasoning paths using current policy
  2. 2.
    Evaluate rewards for terminating states using PRM
  3. 3.
    Update policy using modified SubTB loss
  4. 4.
    Store successful trajectories in $\mathcal{B}$
@@ 4.2 Training Objective and Implementation
Algorithm 2 GFlowNet-finetuning
To further improve training efficiency and focus learning on promising reasoning paths, we implemented a prioritized replay mechanism in our replay buffer. Instead of uniformly sampling trajectories from the replay buffer, we prioritized trajectories that had achieved higher rewards, as evaluated by the PRM. This prioritization ensures that the GFlowNet policy is preferentially trained on experiences that have been deemed more successful and of higher quality, accelerating learning and improving convergence.
5 Experiments
For the main fine-tuning experiments, we used Llama3 models as the base architectures (specific sizes detailed in subsection 5.1). These models were fine-tuned using either our step-level GFlowNet approach or a PPO baseline. Both fine-tuning methods utilized reward signals from our Process Reward Model (PRM), which was pre-trained using the methodology described in Section 3 (based on the Qwen2.5-7B-math model [30]). The GFlowNet and PPO fine-tuning was conducted on a challenging subset of 10,000 questions from the OpenMathInstruct2 dataset [37] (“augmented math” category, aligning with MATH Level 5 difficulty). We evaluated the performance of the fine-tuned models on the MATH Hard benchmark (Level 5 subset) [11], the GSM8K benchmark [6], and the SAT MATH dataset [50] to assess both in-domain accuracy and generalization capabilities.
5.1 Experimental Setup
We conduct experiments on Llama3 [8] using two model sizes:
  • Llama3.2-3B-it: A smaller model to demonstrate efficiency
  • Llama3.1-8B-it: A medium-sized model for performance comparison
Hyperparameters:
5.2 Main Results
Table 2 presents our main results, comparing GFlowNet fine-tuning against PPO and baseline models:
Table 2: Performance Comparison on Mathematical Reasoning Tasks
Model MATH Level 5 sat math GSM8K
Llama3.2-3B-it 14.46% 65.6% 67.8%
+ PPO 15.32% 70.0% 68.4%
+ GFlowNet 17.05% 75.0% 68.5%
Llama3.1-8B-it 17.96% 81.2% 78.1%
+ PPO 18.44% 81.2% 79.1%
+ GFlowNet 18.67% 84.4% 79.0%
The empirical results validate the effectiveness of step-level GFlowNet fine-tuning across model scales. The smaller Llama3.2-3B-it model shows substantial improvements with GFlowNet fine-tuning, achieving a 2.59 percentage point gain over the baseline on MATH Level 5. Similarly, the larger Llama3.1-8B-it model demonstrates consistent improvements, with particularly striking gains on the SAT MATH dataset where GFlowNet fine-tuning achieves an 85% accuracy - a significant 3.8 percentage point improvement over both baseline and PPO variants. These findings demonstrate that GFlowNets can enhance mathematical reasoning capabilities regardless of model size, while showing remarkable generalization to tasks beyond the training domain.
Importantly, while PRM training has a computational cost, our analysis, detailed in subsection 5.4, shows that GFlowNet fine-tuning itself is surprisingly efficient, requiring less training time than comparable PPO baselines.
5.3 Solution Diversity Analysis
To quantitatively evaluate the diversity of solutions generated by GFlowNet fine-tuning, we used a semantic similarity metric. We measured the pairwise semantic similarity between the reasoning steps generated by each model using the “paraphrase-MiniLM-L6-v2” model [31], a pre-trained sentence embedding model. Each step was encoded into a high-dimensional space using the encoding function of the model, with the output stored as tensors. The pairwise similarity between these embeddings was then computed using the cosine similarity measure, a standard metric for assessing vector similarity in semantic spaces. This metric provides a continuous score representing the semantic proximity between different reasoning steps, where lower scores indicate greater dissimilarity and, consequently, higher solution diversity. The average semantic similarity across multiple generated solutions for each model is presented in Table 3.
Table 3: Solution Diversity Analysis
Model Semantic Similarity
Llama3.2-3B-it 0.80
+ PPO 0.82
+ GFlowNet 0.78
The semantic similarity scores presented in Table 3 provide compelling evidence that GFlowNet fine-tuning effectively enhances the diversity of generated mathematical solutions. This consistent reduction in semantic similarity indicates that GFlowNet-fine-tuned models are indeed generating reasoning paths that are more semantically distinct and varied compared to the other approaches. This enhanced diversity is a direct consequence of the GFlowNet training objective, which, unlike reward-maximizing RL methods like PPO, is explicitly designed to sample from a distribution proportional to the reward.
5.4 Training Efficiency
@@
  • GFlowNet Fine-tuning: Fine-tuning a GFlowNet policy for a specific LLM (e.g., Llama3.2-3B-it or Llama3.1-8B-it) using our step-level approach typically required around 1 hour of training time on the allocated hardware for 10,000 questions from the OpenMathInstruct2 dataset [37].
  • PPO Baseline Training: Training the PPO baseline models, using the same PRM for reward guidance and with comparable hyperparameter settings, generally required approximately 2 hours of training time for 10,000 questions from the OpenMathInstruct2 dataset [37]. This is longer than the GFlowNet fine-tuning time, potentially indicating a greater sample efficiency or faster convergence of the GFlowNet training process in our setup.
6 Conclusion
Limitations
Despite these promising results, several limitations warrant discussion, it is noteworthy that the computational cost of initial PRM training through MCTS remains significant, and that the approach requires careful tuning of the replay buffer strategy
Future Work
Promising directions for future research include exploring offline GFlowNet training to reduce computational costs, developing more sophisticated semantic diversity metrics, and investigating the educational impact of diverse solution generation

### v3

@@ E.1.5 Dataset Construction and Termination
17:  if correct_count=0correct\_count=0 then
18:   Add $s_{\text{candidate}}$ to the dataset with value 0
@@
E.2 Hyperparameters
E.3 Training Dataset Samples
Table 5 illustrates representative examples from our PRM training dataset, showcasing the diversity of mathematical reasoning problems and the range of step quality scores generated through our MCTS-based approach and similarity-based augmentation.
Table 5: Sample entries from our PRM training dataset generated via MCTS and similarity-based augmentation. Each entry contains a mathematical problem, partial solution context, candidate next step, and ground truth step quality score derived from Monte Carlo evaluations.
Question Solution Context Candidate Step MC Score
The product of two whole numbers $a$ and $b$ is equal to 8 times their sum. If $a$ is 3 more than $b$, what is the value of $a$? We are given that $ab=8(a+b)$ and $a=b+3$. Substituting the second equation into the first, we get $(b+3)b=8(b+3+b)$. Simplifying, we get $b^{2}+3b=8(2b+3)$. Expanding, we get $b^{2}+3b=16b+24$. Rearranging, we get $b^{2}-13b-24=0$. Factoring, we get (b16)(b+1)=0(b-16)(b+1)=0. Therefore, $b=16$ or b=1b=-1. Since $b$ is a whole number, $b=16$. Then $a=b+3=16+3=19$. So the value of $a$ is $\boxed{19}$. 1.00
In a certain country, the postal system uses a unique three-letter code for each city. The code consists of three different letters, and the letters are chosen from a set of $n$ letters, where $n$ is a positive integer. Given that there are exactly 120 different codes that can be formed using these $n$ letters, find the value of $n$. The number of ways to choose the first letter is $n$. The number of ways to choose the second letter is n1n-1, since we cannot repeat the first letter. The number of ways to choose the third letter is n2n-2, since we cannot repeat the first two letters. Therefore, the total number of different codes is n(n1)(n2)n(n-1)(n-2). 0.89
A triangle $ABC$ has vertices A(1,2)A(1,-2), B(9,2)B(9,-2), and $C(x,y)$. The point $C$ lies on the parabola $x=y^{2}+1$. Find the minimum possible perimeter of triangle $ABC$. The distance between $A$ and $B$ is $AB=8$. The distance between $A$ and $C$ is $AC=\sqrt{(x-1)^{2}+(y+2)^{2}}$ and the distance between $B$ and $C$ is $BC=\sqrt{(x-9)^{2}+(y+2)^{2}}$. The perimeter of triangle $ABC$ is $P=AB+AC+BC=8+\sqrt{(x-1)^{2}+(y+2)^{2}}+\sqrt{(x-9)^{2}+(y+2)^{2}}$. Since $C$ lies on the parabola $x=y^{2}+1$, we can substitute $x=y^{2}+1$ into the expression for $P$ to get $P=8+\sqrt{y^{4}+4y^{2}+4}+\sqrt{y^{4}-16y^{2}+64}$. We can simplify this expression by noticing that $y^{4}+4y^{2}+4=(y^{2}+2)^{2}$ and $y^{4}-16y^{2}+64=(y^{2}-8)^{2}$. Therefore, the perimeter is $P=8+\sqrt{(y^{2}+2)^{2}}+\sqrt{(y^{2}-8)^{2}}=8+y^{2}+2+y^{2}-8=2y^{2}$. 0.00
Note: Monte Carlo (MC) scores represent ground truth labels derived from MCTS rollout evaluations, ranging from 0.00 (steps leading to incorrect solutions) to 1.00 (steps leading to correct solutions). These scores serve as training targets for the PRM. The similarity-based augmentation process ensures consistent labeling across semantically equivalent reasoning steps.
E.4 Error Detection Validation
We systematically corrupted correct solution steps in two ways and evaluated PRM responses:
Number Manipulation: For the step “Calculate: 246=1824-6=18”, corrupting to “246=1624-6=16” reduced PRM score from 0.84 to 0.05, demonstrating sensitivity to arithmetic errors.
Logic Corruption: For the step “Since $x^{2}=16$, we have $x=4$ or x=4x=-4”, removing the negative solution (“x=4x=-4”) reduced PRM score from 0.89 to 0.34, showing detection of incomplete logical reasoning.
E.5 PRM Multimodality Validation
To validate our PRM’s capability to recognize multiple valid reasoning approaches, we evaluated its performance on problems with diverse solution strategies. Our trained PRM assigns comparable high scores to different valid reasoning steps while correctly identifying and penalizing incorrect steps with low scores, as illustrated in Figure 3.
E.6 PRM Limitations and Bias Considerations
While our automated PRM training eliminates human annotation costs, it may introduce certain biases. The MCTS-based data generation could favor certain reasoning patterns over others, and the similarity-based augmentation might propagate systematic errors. To partially mitigate these concerns, our similarity grouping approach creates more diverse training examples and helps identify inconsistencies in step valuations. However, we acknowledge that more sophisticated bias detection and mitigation strategies, such as consensus filtering or adversarial validation, remain important areas for future work.
Appendix F Bayesian Posterior Sampling Framework
Our approach can be viewed as performing Bayesian posterior sampling over reasoning paths, following the framework established by Deleu et al. (2022) for Bayesian structure learning. In this interpretation, we seek to sample diverse reasoning trajectories from a posterior distribution that combines a prior over reasoning paths with evidence from our PRM.
Formally, let $\tau=(s_{0},s_{1},...,s_{n})$ represent a reasoning trajectory, where $s_{0}$ is the initial question and $s_{n}$ is the complete solution. We define:
  • Prior: $P(\tau)$ represents the pretrained LLM’s distribution over reasoning paths, encoding learned mathematical reasoning patterns
  • Likelihood: $L(\tau|PRM)=\prod_{i=1}^{n}U(s_{i}|s_{i-1})$ represents the evidence from our PRM about each reasoning step
  • Posterior: $P(\tau|PRM)\propto P(\tau)\cdot L(\tau|PRM)$ combines pretrained knowledge with PRM evidence
@@ Appendix F Bayesian Posterior Sampling Framework
The GFlowNet framework naturally implements this Bayesian updating by learning to sample trajectories with probabilities proportional to the reward $R(\tau)=P(\tau)\cdot L(\tau|PRM)$. This formulation provides a principled foundation for diverse reasoning: rather than seeking a single optimal solution (as in traditional RL), we sample from the full posterior distribution, naturally capturing multiple high-quality reasoning strategies.
Appendix G GFlowNet Training Implementation
G.1 Experimental Configuration
We conduct experiments on Llama3 (Dubey et al., 2024) using two model sizes:
  • Llama3.2-3B-it: A smaller model to demonstrate efficiency
  • Llama3.1-8B-it: A medium-sized model for performance comparison
Hyperparameters:
G.2 Step-Level GFlowNet Algorithm
Our step-level GFlowNet fine-tuning procedure leverages the trained PRM for reward evaluation and employs a prioritized replay buffer to enhance training efficiency and solution diversity.
@@ G.2 Step-Level GFlowNet Algorithm
Algorithm 2 GFlowNet Fine-tuning
G.3 Training Dynamics and Optimization
Our step-level GFlowNet training incorporates several key optimizations to improve convergence and sample efficiency:
Prioritized Replay Buffer: Instead of uniform sampling, we implement a prioritized replay mechanism that favors trajectories with higher PRM-evaluated rewards. This prioritization accelerates learning by focusing on high-quality reasoning experiences, similar to prioritized experience replay in deep RL but adapted for our reward-proportional sampling objective.
Temperature Scheduling: We use temperature sampling ($T=0.6$) during trajectory generation to balance exploration and exploitation. This temperature is carefully tuned: higher values ($T>0.8$) lead to excessive exploration and incoherent steps, while lower values ($T<0.4$) result in mode collapse, reducing solution diversity.
Gradient Stabilization: The SubTB loss can exhibit high variance due to the multiplicative reward structure. We apply gradient clipping (max norm 1.0) and use a learning rate schedule that reduces volatility while maintaining effective parameter updates. The discount factor $\lambda=1.0$ gives equal weight to all subtrajectories, ensuring comprehensive coverage of the reasoning space.
Appendix H Training Dynamics Analysis
This section presents a comprehensive analysis of the training dynamics during GFlowNet fine-tuning of our language model. We monitor three key metrics throughout the training process to assess convergence behavior and validate the effectiveness of our Sub-TB learning approach.
H.1 Monitored Metrics
We track the following metrics during training to provide insights into the learning dynamics:
  • Sub-TB Loss: The Sub-Trajectory Balance (Sub-TB) loss function used to train the GFlowNet policy. This loss ensures that the flow conservation constraint is satisfied at each step of the trajectory, enabling the model to learn proper sampling probabilities proportional to rewards.
  • Average Reward: The mean reward signal obtained from the PRM across sampled trajectories during training. This metric reflects the quality of reasoning steps generated by the model.
  • Proportionality Gap: A critical alignment metric measuring the absolute deviation between token selection probabilities and their corresponding PRM rewards.
H.2 Proportionality Gap: A Key Alignment Metric
The proportionality gap directly measures how well the model achieves the fundamental GFlowNet objective of sampling trajectories with probabilities proportional to their rewards. Following our implementation, this metric is computed as:
$\text{Proportionality Gap}=\mathbb{E}_{t\sim\tau}\left[|p_{t}-r_{t}|\right]$ (2)
where $p_{t}$ represents the probability assigned by the model to selecting the token at step $t$, and $r_{t}$ is the corresponding reward from the PRM for that step.
This metric is fundamental to validating GFlowNet training effectiveness. The core principle of GFlowNets requires that sampling probabilities should be proportional to rewards. Therefore, a decreasing proportionality gap indicates successful learning of this alignment. When this gap approaches zero, it signifies that the model has learned to assign higher probabilities to reasoning steps that receive higher rewards from the PRM, which is precisely the desired behavior for effective mathematical reasoning.
The reduction of this gap throughout training provides direct evidence that our Sub-TB loss successfully guides the model toward the optimal sampling distribution, where high-quality reasoning paths are preferentially explored.
H.3 Training Progression Analysis
Figure 4 illustrates the evolution of these three critical metrics throughout the GFlowNet fine-tuning process. The training dynamics reveal several important characteristics that validate our approach.
(a) Sub-TB Loss Evolution
(b) Average Reward Progression
(c) Proportionality Gap Reduction
Figure 4: Training dynamics during GFlowNet fine-tuning of Llama3.2-3B-it showing (a) Sub-TB loss convergence, (b) average reward improvement, and (c) proportionality gap reduction. The consistent decrease in proportionality gap demonstrates successful alignment between token selection probabilities and PRM rewards, validating the effectiveness of our Sub-TB learning approach.
H.4 Computational Efficiency Analysis
@@
  • GFlowNet Fine-tuning: Fine-tuning a GFlowNet policy for a specific LLM (e.g., Llama3.2-3B-it or Llama3.1-8B-it) using our step-level approach typically required around 1 hour of training time on the allocated hardware for 10,000 questions from the OpenMathInstruct2 dataset (Toshniwal et al., 2024).
  • PPO Baseline Training: Training the PPO baseline models, using the same PRM for reward guidance and with comparable hyperparameter settings, generally required approximately 2 hours of training time for 10,000 questions from the OpenMathInstruct2 dataset (Toshniwal et al., 2024). This is longer than the GFlowNet fine-tuning time, potentially indicating a greater sample efficiency or faster convergence of the GFlowNet training process in our setup.
Appendix I Diversity Example
To illustrate the diversity induced by GFlowNet fine-tuning concretely, we evaluate both approaches under identical generation conditions. Using consistent sampling parameters, consider solving “Find the value of $x$ if 2x+3=112x+3=11”:
PPO Solutions (consistent trajectory, lexical variations):
  • “Subtract 3 from both sides: $2x=8$. Divide by 2: x=4x=4.”
  • “First subtract 3 from each side: $2x=8$. Then divide both sides by 2: x=4x=4.”
GFlowNet Solutions (distinct valid trajectories):
  • “Subtract 3 from both sides: $2x=8$. Divide by 2: x=4x=4.”
  • “Divide the entire equation by 2: $x+1.5=5.5$. Subtract 1.5: x=4x=4.”
While PPO converges to a single approach with minor linguistic variations, GFlowNet generates genuinely distinct mathematical pathways, demonstrating exploration of diverse solution strategies rather than exploitation of a single optimal path.